# Mapping global offshore wind resource: wake losses, optimisation potential and climate effects

### v1

@@ Abstract
In this work, we assess global offshore wind energy resources, wake-induced losses, array layout optimisation potential and climate change impacts. We first map global offshore ambient wind resource from reanalysis data. We estimate wake-induced losses using an engineering wake model, revealing that locations with low (high) resource typically experience larger (smaller) percentage losses. However, we further find that the specific wind speed distribution is important, with narrower distributions generally leading to greater losses. This is due to the overlap between the wind speed distribution and the high-sensitivity region of the turbine thrust and power curves. Broadly, this leads to much stronger wake-induced losses in the tropics (which experience the trade winds) than mid-latitudes. However, the tropics also experience a narrower wind direction distribution; our results demonstrate that this leads to greater potential for mitigation of wake effects via layout optimisation. Finally, we assess projected changes in wind resource and wake losses due to climate change under a high-emission scenario. Many regions are projected to decrease in ambient wind resources, and furthermore these regions will typically experience greater wake-induced losses, exacerbating the climate impact. These results highlight the different challenges and opportunities associated with exploiting offshore wind resources across the globe.
@@ 1 Introduction
A number of popular tools exist that are dedicated to mapping wind resources. These include so-called ‘wind atlas’ products, such as the New European Wind Atlas (Hahmann et al., 2020, Dörenkämper et al., 2020) and Global Wind Atlas (Davis et al., 2023). Both are based on ERA-5 hindcast data (Hersbach et al., 2020), dynamically downscaled using the Weather Research and Forecasting (WRF) model (Powers et al., 2017), with ‘microscale’ modelling via the WAsP software suite (Mortensen et al., 2001). Reanalysis products are also used directly for wind resource mapping (e.g. Soares et al., 2020, Gruber et al., 2022). Either reanalyses or wind atlases are commonly used as inputs into energy systems models (ESMs), which are in turn used to inform policy decisions regarding future energy mix and infrastructure. One popular tool providing ESM wind power inputs is the Renewables.Ninja project (Staffell and Pfenninger, 2016), which is based on bias-corrected reanalysis data. In the climate context, wind resource mapping is commonly undertaken using either global projection model outputs such as from CMIP5 (e.g. Reyers et al., 2016, Zheng et al., 2019) or CMIP6 (e.g. Ibarra-Berastegui et al., 2023), or using downscaled climate products such as from the CORDEX project (e.g. Li et al., 2020, Molina et al., 2022, Akperov et al., 2023), or other downscaling methods (e.g. Zhang and Li, 2021, Fernández-Alvarez et al., 2023).
An important aspect to assessing potential wind resource extraction is power loss due to turbine wake effects (Jensen, 1983). The extraction of kinetic energy by a wind turbine reduces the wind speed immediately downstream, impacting the power output from any downstream turbines within the same farm (so-called ‘intra-farm’ or ‘internal’ wakes). This effect also arises at larger spatial scales, where an upstream wind farm can reduce the power output of a downstream farm (Lundquist et al., 2019, Schneemann et al., 2020, Cañadillas et al., 2022), resulting in so-called ‘inter-farm’ or ‘cluster’ wakes. In the intra-farm case, average wake effects will depend on the turbine type and the layout of the wind farm, and in the inter-farm case on the size, density, capacity and relative positions of the farms. However, in both cases, another important factor is the local distribution of wind speed and direction. The wind speed is important since wake effects can only impact power output while the wind speed is greater than the turbine’s cut-in speed, but not significantly greater than the rated speed. The wind direction is important because a given pair of turbines or farms only interact when the wind direction is (broadly) aligned with their relative positions.
@@
However, despite the breadth of literature on wake effects, their incorporation into resource assessment tools or studies is limited. In the climate context, Hahmann et al. (2022) use an engineering wake model to include wake-induced losses of a proposed wind farm cluster, within a study of the climate impact on wind energy resources in the North Sea. Warder and Piggott (2024) study the evolution of wake effects (on various spatial scales) with climate change, based on a near-future set of wind farms in the German Bight, finding statistically significant changes in wake effects during the summer months. In the reanalysis context, and in particular for models used as inputs into ESMs, wake effects may be implicitly accounted for via bias correction methods based on matching models to observed capacity factors (Staffell and Pfenninger, 2016, Benmoufok et al., 2024). However, in this case, variation in wake effects either spatially or due to differing wind farm designs would be neglected. Alternatively, wake effects are sometimes explicitly accounted for, e.g. via a fixed ‘array efficiency’ factor (Bosch et al., 2018, 2019), or using pre-computed losses for idealised farms of various sizes (Glaum et al., 2024), such as those found in Volker et al. (2017), Energiewende et al. (2020), Gea-Bermúdez et al. (2022). In reality, however, wake effects vary depending on the specific wind climate experienced at a given location, as well as with climate change (Devis et al., 2018). Furthermore, the potential to mitigate these losses via array layout optimisation will vary with the wind direction distribution, with narrower distributions likely to offer greater optimisation potential. These factors will have implications for the transfer of knowledge and experience from regions such as Europe, where the offshore wind industry is relatively mature, to emerging markets elsewhere in the world, where the challenges and opportunities associated with wake effects may differ significantly.
With a focus on global spatial variation, the objectives of the present study are to investigate
@@
the ambient offshore wind resource;
@@
the power losses due to intra-farm wake effects for a given wind farm;
@@
the layout optimisation potential, which we define as the percentage improvement in wind farm power which can be achieved via layout optimisation, compared with a simple grid layout;
@@
the impact of climate change on both ambient offshore wind resource, and intra-farm wake effects.
The remainder of the paper is structured as follows. Section 2 presents the data and methods used within this study, including reanalysis data, climate projection data, and wake modelling and layout optimisation approaches. Corresponding results are presented and discussed in section 3. Finally, we draw conclusions in section 4.
@@ 2.1.1 ERA-5
We utilise reanalysis data from ERA-5 (Hersbach et al., 2020), for the historical period 1995–2014. We use the 100 and wind components, from which we derive the wind speed and direction. The data has a spatial resolution of 0.25, and hourly temporal resolution. We download data from the Copernicus Climate Data Store at a 6-hourly resolution. From this, we derive the wind rose for each grid cell, based on wind speed bins from 0 to 25 in steps of 0.5 , and directional bins of 10. The use of a wind rose to represent the wind distribution at each location reduces the total volume of data required to represent global wind climate at high resolution.
@@ 2.1.2 CMIP6
The future climate scenario we assess in this work is the SSP585 scenario, from which we use the period 2081–2100, using CMIP6 projections. This high-emissions scenario corresponds to intensified fossil fuel exploitation, leading to radiative forcing of 8.5 by the year 2100, and lies at the upper end of the scenarios considered in the literature. We choose this scenario on the basis that it will produce responses of the greatest magnitude, making it easiest to identify robust trends. We compare this future scenario against corresponding CMIP6 data from the ‘historical’ scenario, using the period 1995–2014.
@@
The specific CMIP6 model outputs we select for this study are based on the availability of both the ‘historical’ and SSP585 scenarios, a nominal resolution of 100 or finer, and surface wind speeds with daily frequency. Data was searched and downloaded via the Earth System Grid Federation (ESGF) search tool (https://esgf-ui.ceda.ac.uk/cog/search/cmip6-ceda/). The resulting 10 models are detailed in table 1.
@@
Each of these models provides the ‘surface’ wind speed, at a nominal height of 10 . Since we are interested in wind speeds at hub heights, we extrapolate these data to 100 . For this, we use a power law with an exponent of 0.06, following Carvalho et al. (2021), who chose this value based on typical vertical profiles from ERA-5 data, which is available at both 10 and 100 . The resulting data is converted to a wind rose map (consistent with our treatment of the ERA-5 data described above), then regridded by bilinear interpolation to match the ERA-5 grid.
@@
We further bias-correct the climate projection data against the ERA-5 data. Warder and Piggott (2024) compared several common methods for bias correction, with a quantile mapping method based on fitted Weibull distributions found to perform best. However, since wind speeds are not necessarily well described by Weibull distributions at all locations globally, in this study we employ an empirical quantile mapping method, which was also found by Warder and Piggott (2024) to perform well. The bias correction is derived based on the climate dataset over the period that coincides with the historical ERA-5 data, using the latter as ‘ground truth’. Denoting the empirical cumulative density function (CDF) of the ERA-5 data at a particular grid cell and for a particular wind direction as , and the corresponding CDF from the ‘historical’ climate dataset as , the corrected wind speed is given as a function of the raw wind speed as
@@
Since our approach is based on wind roses, the projected climate data is represented as a discrete probability density function (PDF) at a particular set of speeds. Denoting the raw PDF as and the corrected PDF as , we assume that
@@
The raw PDF therefore directly provides the corrected PDF, but evaluated at the ‘corrected’ speeds given by equation (1). To evaluate the corrected PDF on the original set of wind speeds, we linearly interpolate within the values obtained by equation (2).
@@
2.1.3 Other data
We use global data for marine exclusive economic zones (EEZs) from Flanders Marine Institute (2020), in order to restrict our analysis to plausible offshore regions, and to aggregate results for different countries. However, we leave the inclusion of other practical constraints to wind farm deployment, such as bathymetry, shipping lanes etc, to future work. This study uses a resolution of 0.25, which would not be sufficient to adequately resolve all such features.
@@ 2.2 Mapping wake effects
This study seeks to investigate the influence of wind speed distribution on the magnitude of intra-farm wake effects. Our approach is based on computing wind roses at each grid cell whose centroid is within the EEZs. For each such grid cell, we compute two quantities:
@@
the ‘ambient’ capacity factor which would be produced by a single turbine within that cell, or equivalently the capacity factor of a wind farm, neglecting all wake effects;
@@
The reference wind turbine we use within this work is the 10 reference turbine from IEA Task 37 (Bortolotti et al., 2019), for which power and thrust curve data is available. Neglecting conversion efficiency, the rated (mechanical) power of this turbine is 10.6 . The turbine has a hub height of 119 and a rotor diameter of 198 . As our ‘standard’ farm, we take a 10 by 10 square array of such wind turbines, with a spacing of 5 rotor diameters. This farm has a total installed capacity of 1.06 . The choice of wind turbine model, turbine separation and total capacity were selected in order to be broadly representative of current and near-future offshore wind farms.
The wind rose at a given grid cell is a discrete probability density function, represented by a two-dimensional array which we denote , where and denote a particular wind speed and direction, respectively. The first quantity above, the ambient capacity factor, does not depend on the wind direction. We therefore introduce the marginal wind speed distribution, given by
@@
For the second quantity above, the waked capacity factor, our initial focus is on the effect of the wind speed distribution. We therefore again take the marginal wind speed distribution, and for each wind speed, we estimate the power produced by the farm, averaged over all possible rotations of the farm. This is equivalent to assuming that all wind directions are equally likely at a given location. This also has the effect of reducing the sensitivity of the waked power with respect to the specific turbine array layout we have chosen, and the local prevailing wind. I.e., here we wish to exclude variations in wake-induced losses which arise from the alignment (or otherwise) of the principal directions of the farm’s grid layout with the prevailing wind.
To estimate the power output of the wind farm, at a particular 100 wind speed and subject to a given rotation (or equivalently a given wind direction) , we use the engineering wake model of Bastankhah and Porté-Agel (2014), as implemented within the PyWake software suite (Pedersen et al., 2019). Specifically, we use the predefined ‘IEA37SimpleBastankhahGaussian’ wind farm model available within PyWake. This model assumes a fixed wake expansion parameter of , and therefore neglects the influence of turbulence (or any other factors) on wake expansion (Peña and Rathmann, 2014). We further neglect the influence of air density on power production or wakes. These assumptions enable us to isolate the influence of wind speed distributions on power and wakes, and we leave further exploration of their impact to future work.
The simulated power output of our ‘standard’ farm for speed and rotation (or wind direction) is denoted . We then compute the mean power for each wind speed as
where is the number of wind directions, for which we choose 36, or bins of 10. This mean power is then combined with the marginal wind speed distribution , to produce the final ‘waked’ capacity factor, given by
where is the total number of wind turbines. We emphasise that by computing the capacity factor in this way, we are discarding all information about the wind direction distribution, and isolating the influence of the wind speed distribution. This allows for greater insight into the influence of wind speed distribution on power and wake effects, and reduces the sensitivity of the results to our choice of wind turbine layout within our ‘standard’ farm.
As well we the raw and values, we also present results for the capacity factor lost to wakes, given by

### v2

# Mapping global offshore wind wake losses, layout optimisation potential, and climate change effects

@@ Abstract
This study assesses global offshore wind energy resources, wake-induced losses, array layout optimisation potential and climate change impacts. Global offshore ambient potential is first mapped based on reanalysis data. Wake-induced losses are then estimated using an engineering wake model, revealing that locations with low (high) resource typically experience larger (smaller) percentage losses. However, the specific wind speed distribution is found to be important, with narrower distributions generally leading to greater losses. This is due to the overlap between the wind speed distribution and the high-sensitivity region of the turbine thrust and power curves. Broadly, this leads to much stronger wake-induced losses in the tropics (which experience the trade winds) than mid-latitudes. However, the tropics also experience a narrower wind direction distribution; the results of this study demonstrate that this leads to greater potential for mitigation of wake effects via layout optimisation. Finally, projected changes in wind potential and wake losses due to climate change under a high-emission scenario are assessed. Many regions are projected to decrease in ambient wind resources, and furthermore these regions will typically experience greater wake-induced losses, exacerbating the climate change impact. These results highlight the different challenges and opportunities associated with exploiting offshore wind resources across the globe.
@@ 1 Introduction
A number of popular tools exist that are dedicated to mapping wind resources. These include so-called ‘wind atlas’ products, such as the New European Wind Atlas (Hahmann et al., 2020, Dörenkämper et al., 2020) and Global Wind Atlas (Davis et al., 2023). Both are based on ERA-5 hindcast data (Hersbach et al., 2020), dynamically downscaled using the Weather Research and Forecasting (WRF) model (Powers et al., 2017), with ‘microscale’ modelling via the WAsP software suite (Mortensen et al., 2001). Reanalysis products are also used directly for wind resource mapping (e.g. Soares et al., 2020, Gruber et al., 2022). Either reanalyses or wind atlases are commonly used as inputs into energy systems models (ESMs), which are in turn used to inform policy decisions regarding future energy mix and infrastructure. One popular tool providing ESM wind power inputs is the Renewables.Ninja project (Staffell and Pfenninger, 2016), which is based on bias-corrected reanalysis data. In the climate change context, wind resource mapping is commonly undertaken using either global projection model outputs such as from CMIP5 (e.g. Reyers et al., 2016, Zheng et al., 2019) or CMIP6 (e.g. Ibarra-Berastegui et al., 2023), or using downscaled climate products such as from the CORDEX project (e.g. Li et al., 2020, Molina et al., 2022, Akperov et al., 2023), or other downscaling methods (e.g. Zhang and Li, 2021, Fernández-Alvarez et al., 2023).
An important aspect to assessing potential wind resource extraction is power loss due to turbine wake effects (Jensen, 1983). The extraction of kinetic energy by a wind turbine reduces the wind speed immediately downstream, impacting the power output from any downstream turbines within the same farm (so-called ‘intra-farm’ or ‘internal’ wakes). This effect also arises at larger spatial scales, where an upstream wind farm can reduce the power output of a downstream farm (Lundquist et al., 2019, Schneemann et al., 2020, Cañadillas et al., 2022), resulting in so-called ‘inter-farm’ or ‘cluster’ wakes. In the intra-farm case, average wake effects will depend on the turbine type and the layout of the wind farm, and in the inter-farm case on the size, density, capacity and relative positions of the farms (Wang et al., 2023b). However, in both cases, another important factor is the local distribution of wind speed and direction. The wind speed is important since wake effects can only impact power output while the wind speed is greater than the turbine’s cut-in speed, but not significantly greater than the rated speed. The wind direction is important because a given pair of turbines or farms only interact when the wind direction is (broadly) aligned with their relative positions.
@@
However, despite the breadth of literature on wake effects, their incorporation into resource assessment tools or studies is limited. In the climate change context, Hahmann et al. (2022) use an engineering wake model to include wake-induced losses of a proposed wind farm cluster, within a study of the climate impact on wind energy resources in the North Sea. Warder and Piggott (2025) study the evolution of wake effects (on various spatial scales) with climate change, based on a near-future set of wind farms in the German Bight, finding statistically significant changes in wake effects during the summer months. In the reanalysis context, and in particular for models used as inputs into ESMs, wake effects may be implicitly accounted for via bias correction methods based on matching models to observed capacity factors (Staffell and Pfenninger, 2016, Benmoufok et al., 2024). However, in this case, variation in wake effects either spatially or due to differing wind farm designs would be neglected. Alternatively, wake effects are sometimes explicitly accounted for, e.g. via a fixed ‘array efficiency’ factor (Bosch et al., 2018, 2019), or using pre-computed losses for idealised farms of various sizes (Glaum et al., 2024), such as those found in Volker et al. (2017), Energiewende et al. (2020), Gea-Bermúdez et al. (2022). In reality, however, wake effects vary depending on the specific wind climate experienced at a given location, as well as with climate change (Devis et al., 2018). Furthermore, the potential to mitigate these losses via array layout optimisation will vary with the wind direction distribution, with narrower distributions likely to offer greater optimisation potential. These factors will have implications for the transfer of knowledge and experience from regions such as Europe, where the offshore wind industry is relatively mature, to emerging markets elsewhere in the world, where the challenges and opportunities associated with wake effects may differ significantly.
Two key gaps are identified in the above literature, which are directly addressed in the present study. Firstly, to the authors’ knowledge, there has been no globally consistent assessment of the impact of wake effects on the ability to extract wind power, and the scope to mitigate these losses via array layout optimisation; these factors are essential to the transfer of knowledge to emerging offshore wind markets. Secondly, there has been no global assessment of projected changes in wake-induced losses due to climate change; this has significant implications for long-term planning and policy makers. Specifically, with a focus on global spatial variation, the objectives of this study are to investigate
@@
ambient offshore wind potential;
@@
the power losses due to intra-farm wake effects for a selected wind farm;
@@
the layout optimisation potential, defined as the percentage improvement in wind farm power which can be achieved via layout optimisation, compared with a simple grid layout;
@@
the projected impact of climate change on both ambient offshore wind potential, and intra-farm wake effects, under the SSP585 scenario.
The remainder of the paper is structured as follows. Section 2 presents the data and methods used within this study, including reanalysis data, climate projection data, and wake modelling and layout optimisation approaches. Corresponding results are presented and discussed in section 3. Finally, conclusions are drawn in section 4.
@@ 2 Methods
The workflow employed within this study is summarised in figure 1. Resource data is taken from ERA-5 reanalysis, as well as CMIP6 climate projections. In the latter case, bias correction is performed using ERA-5 data as ground truth. Along with data describing global marine exclusive economic zones (EEZs), and an assumed wind farm layout and turbine power curve, these resource datasets are used to generate global maps of both ambient wind potential, and wake-induced losses. The wake modelling utilises an engineering wake model within the PyWake software suite. Wind turbine layout optimisation is subsequently performed using the TopFarm package, in order to map the ‘optimisation potential’ of wind farms across marine EEZs. Further detail on each of the components of this workflow is given in the following subsections.
@@ 2.1.1 ERA-5
This study utilises reanalysis data from ERA-5 (Hersbach et al., 2020), for the historical period 1995–2014. The wind speed and direction are derived from the 100 and wind components from this ERA-5 data. The data has a spatial resolution of 0.25, and hourly temporal resolution. The data is downloaded from the Copernicus Climate Data Store at a 6-hourly resolution. From this, the wind rose is derived for each grid cell, based on wind speed bins from 0 to 25 in steps of 0.5 , and directional bins of 10. The use of a wind rose to represent the wind distribution at each location reduces the total volume of data required to represent global wind climate at high resolution.
@@ 2.1.2 CMIP6
The future climate scenario assessed in this work is derived from CMIP6 projections, and specifically the SSP585 scenario, for the period 2081–2100. This high-emissions scenario corresponds to intensified fossil fuel exploitation, leading to radiative forcing of 8.5 by the year 2100, and lies at the upper end of the scenarios considered in the literature. This scenario is chosen on the basis that it will produce responses of the greatest magnitude, making it easiest to identify robust trends. This future scenario is compared against corresponding CMIP6 data from the ‘historical’ scenario, using the period 1995–2014.
@@
The specific CMIP6 model outputs selected for this study are based on the availability of both the ‘historical’ and SSP585 scenarios, a nominal resolution of 100 or finer, and surface wind speeds with daily frequency. While higher spatial and temporal resolution is available for some models, these selection criteria were chosen to ensure that a sufficient range of models was available from which to form an ensemble. Data was searched and downloaded via the Earth System Grid Federation (ESGF) search tool (https://esgf-ui.ceda.ac.uk/cog/search/cmip6-ceda/). The resulting 10 models are detailed in table 1.
@@
Each of these models provides the ‘surface’ wind speed, at a nominal height of 10 . Since power estimation requires wind speeds at hub heights, these 10 data are extrapolated to 100 , using a power law with an exponent of 0.06, following Carvalho et al. (2021), who chose this value based on typical vertical profiles from ERA-5 data, which is available at both 10 and 100 . The resulting data is converted to a wind rose map (consistent with the treatment of the ERA-5 data described above), then regridded by bilinear interpolation to match the ERA-5 grid.
@@
The climate projection data is further bias-corrected against the ERA-5 data. This bias correction is performed separately for each grid cell. Warder and Piggott (2025) compared several common methods for bias correction, with a quantile mapping method based on fitted Weibull distributions found to perform best. However, since wind speeds are not necessarily well described by Weibull distributions at all locations globally, this study employs an empirical quantile mapping method, which was also found by Warder and Piggott (2025) to perform well. The use of a quantile mapping method, rather than a simpler rescaling or offset approach, is also well suited here since the climate datasets may underestimate wind speed variability due to the use of daily mean data. A quantile mapping approach is sufficient to address this issue, since it can amend the shape of the wind speed probability distribution.
The bias correction is derived based on the climate dataset over the period that coincides with the historical ERA-5 data, using the latter as ‘ground truth’. Denoting the empirical cumulative density function (CDF) of the ERA-5 data at a particular grid cell and for a particular wind direction as , and the corresponding CDF from the ‘historical’ climate dataset as , the corrected wind speed is given as a function of the raw wind speed as
@@
Since the approach of this study is based on wind roses, the projected climate data is represented as a discrete probability density function (PDF) at a particular set of speeds. Denoting the raw PDF as and the corrected PDF as , we assume that
@@
The raw PDF therefore directly provides the corrected PDF, but evaluated at the ‘corrected’ speeds given by equation (1). Linear interpolation is then used to evaluate the corrected PDF on the original set of wind speeds.
@@
2.1.3 Marine exclusive economic zones
This study uses global data for marine exclusive economic zones (EEZs) from Flanders Marine Institute (2020), in order to restrict the analysis to plausible offshore regions, and to aggregate results for different countries. However, the inclusion of other practical constraints to wind farm deployment, such as bathymetry, shipping lanes etc, is left to future work. This study uses a resolution of 0.25, which would not be sufficient to adequately resolve all such features.
@@ 2.2 Mapping wake effects
This study seeks to investigate the influence of wind speed distribution on the magnitude of intra-farm wake effects. The approach is based on computing wind roses at each grid cell whose centroid is within the EEZs. For each such grid cell, two quantities are computed:
@@
the ‘ambient’ capacity factor, or ambient potential, which would be produced by a single turbine within that cell, or equivalently the capacity factor of a wind farm, neglecting all wake effects;
@@
The reference wind turbine used within this work is the 10 reference turbine from IEA Task 37 (Bortolotti et al., 2019), for which power and thrust curve data is available. Neglecting conversion efficiency, the rated (mechanical) power of this turbine is 10.6 . The turbine has a hub height of 119 and a rotor diameter of 198 . As a ‘standard’ farm, this study takes a 10 by 10 square array of such wind turbines, with a spacing of 5 rotor diameters. This farm has a total installed capacity of 1.06 . The choice of wind turbine model, turbine separation and total capacity were selected in order to be broadly representative of current and near-future offshore wind farms.
The wind rose at a given grid cell is a discrete probability density function, represented by a two-dimensional array which can be denoted , where and denote a particular wind speed and direction, respectively. The first quantity above, the ambient capacity factor, does not depend on the wind direction. We therefore introduce the marginal wind speed distribution, given by
@@
For the second quantity above, the waked capacity factor, the initial focus of this study is on the effect of the wind speed distribution. This can be achieved by taking the marginal wind speed distribution, and for each wind speed, estimating the power produced by the farm, averaged over all possible rotations of the farm. This is equivalent to assuming that all wind directions are equally likely at a given location. This also has the effect of reducing the sensitivity of the waked power with respect to the specific turbine array layout chosen, and the local prevailing wind. I.e., the aim is to exclude variations in wake-induced losses which arise from the alignment (or otherwise) of the principal directions of the farm’s grid layout with the prevailing wind.
Estimation of the power output of the wind farm, at a particular 100 wind speed and subject to a given rotation (or equivalently a given wind direction) , is simulated using the engineering wake model of Bastankhah and Porté-Agel (2014), as implemented within the PyWake software suite (Pedersen et al., 2019). Specifically, the predefined ‘IEA37SimpleBastankhahGaussian’ wind farm model available within PyWake is used. A fixed wake expansion parameter of is assumed, neglecting the influence of turbulence (or any other factors) on wake expansion. The influence of air density on power production or wakes is neglected. These assumptions enable the isolation of the influence of wind speed distributions on power and wakes, and further exploration of the impact of air density is left to future work. Turbine interactions are computed by starting with the most upwind turbine and iterating in the downstream direction. Squared sum superposition is used to estimate the effect of overlapping wakes. For further detail regarding the implementation of the wake model, the reader is referred to the PyWake documentation (DTU Wind Energy, 2018). This modelling workflow constitutes a commonly-used approach, and the wake model of Bastankhah and Porté-Agel (2014) has been previously validated against a variety of observation data including LES simulations (Bastankhah and Porté-Agel, 2014), lidar (Carbajo Fuertes et al., 2018) and operational data (Rodrigo et al., 2020).
The simulated power output of our ‘standard’ farm for speed and rotation (or wind direction) is denoted . The mean power for each wind speed can then be computed as
where is the number of wind directions, chosen as 36, or equivalently using bins of 10. This mean power is then combined with the marginal wind speed distribution , to produce the final ‘waked’ capacity factor, given by
where is the total number of wind turbines. It should be emphasised that by computing the capacity factor in this way, all information about the wind direction distribution is discarded, isolating the influence of the wind speed distribution. This allows for greater insight into the influence of wind speed distribution on power and wake effects, and reduces the sensitivity of the results to the choice of wind turbine layout within the ‘standard’ farm.
As well as the raw and values, the capacity factor lost to wakes is also presented, given by